D Question 22 2 pts This question is not specific to any previous case: If an individual weighs 175 pounds and typically losses 6 lbs of weight during a 3-hour practice, how much fluid should they consume during exercise if they want to limit thier fluid losses in future workouts to no more than 2%. O Approximately 12 Cups O 16-18 Cups O No less than 7 Cups O No less than 3.5 Cups This is the information provided
Answer #1

The weight of water can be considered as 1g per ml or 2.2 pounds per 1 liter.

This weight of water can be measured as equivalent to 2 cups of liquid lost per one pound of weight loss. As some water is also lost through urinary excretion, 3 cups of fluid loss can be attributed to one pound of weight loss. To replace the lost fluid from the body the amount of water to be taken during a 3-hour exercise is calculated as below.

If 6 pounds of weight is lostà 6*3cups = 18 cups of fluid is lost.

The percent fluid loss for 6 pounds is 600/175 = 3.42%

If the fluid is taken as 1.5 cups for every 15 mints; for 3 hours practice, it becomes 18 cups of water.

So, the above calculation of fluid intake is for 3.42% of fluid loss.

If the fluid loss is not more than 2% then the number of lost pounds of weight should not be more than 3.5pounds. Then the fluid that has to be taken in for replacing the lost weight should be less than 10.5 cups.

3.5*3cups = 10.5 cups.

Among the answers available,

The correct answer would be no less than 7 cups.